# Unveiling the inconsistency of the Proca theory with non-minimal coupling to gravity

### v3

@@ 6 Conclusion
In this work, we have studied Proca theory, non-minimally coupled to gravity. The cosmological aspects of this theory have recently drawn a lot of attention, making it an intriguing possibility to underline the gravitational production of dark matter and a possible candidate for generating primordial magnetic fields and inflationary candidates. By studying this theory in flat space-time, with metric perturbations also taken into account, we have found a somewhat surprising result. On the one hand, the interaction of the vector field with the Ricci scalar results in the strong coupling of the longitudinal modes. This is relatively natural to expect, as one obtains self-interaction of the longitudinal modes at the quartic order. Beyond it, the vector and tensor modes remain weakly coupled, up to the Planck length, due to the amplitude of quantum fluctuations for the tensor modes. On the other hand, the coupling of the vector modes with the Ricci tensor indicates that both tensor and longitudinal modes become strongly coupled at the same Vainsthein scale. This, however, is an unphysical result – it would imply that if one allows for such non-minimal coupling in nature, the gravitational waves would quickly lose their linear propagator. Moreover, a recent study has found a new class of the negative frequency modes in curved space-time, known as the runaway modes [26]. In this work, we have shown that its origin lies precisely in the coupling of the vector field to the Ricci tensor, which is problematic even in the flat background.
@@
Nevertheless, by carefully analyzing the theory at this strong coupling scale, we have found that the tensor modes have to be transformed. A similar scenario was found also in massive Yang-Mills theory, with mass added by hand, where with standard decomposition, the transverse modes were becoming strongly coupled. However, by correcting their definition, it was shown that ultimately, these modes remain weakly coupled at all scales.
In the case of non-minimally coupled Proca theory, we have found that the introduction of corresponding disformal coupling in the theory simultaneously removes the inconsistent interaction by introducing the necessary coupling. This leads us to the self-interaction of the longitudinal modes at the quartic order, becoming a necessary step to make the theory safe. Notably, the same transformation removes the problematic runaway modes, while at the same time generating further self-interactions of the vector field. So it will be interesting to study how the predictions made for the original theory change once this is taken into account, if the runaway could re-appear once the theory is quantized, and to explore further aspects of the strong coupling in curved space-time for this type of theory. In particular, it would be intriguing to see if such a possibility could resolve the well-known ghost, that appears for $M_{T}^{2}<0$.
One potential curiosity might be to investigate the theory in the presence of external matter. For this, one has two possibilities – either to first couple the matter with the old metric, or couple it only after the disformal transformation has been performed. Intuitively, one can notice that even in the kinetic term the scalar will generate quartic interactions in this case. In contrast, if the external matter is added aposteriori, the possibility of the scalar field feeling the interaction could vanish more safely. While it would be interesting to check this analysis for both scenarios, a realistic perspective is that the matter should be minimally coupled to the transformed metric, as in this frame tensor modes are weakly coupled, which is in an agreement with what we observe today. We will leave further aspects of this investigation for the upcoming works. Finally, one should note that while in Proca theory, the longitudinal modes are the source of the singularities in mass within the perturbative regime, in the Kalb-Ramond field, this role is taken by the vector modes [50]. Therefore, it will be interesting to also extend this analysis to this case, and study its consequences.

### v4

@@ 5 Resolving the Runaway Modes
One should note that if (63) is taken as a field redefinition, it may happen that the instability persists in the theory, through the coupling with matter. This will certainly be interesting to check in future works. In this work, however, one should stress that (63) in both section 4 and this section is taken as a definition of the disformal frame, where matter should be minimally coupled to gravity. This is similar to the choice of either the Jordan or the Einstein frame in f(R) theories of gravity for the frame in which the matter is minimally coupled to gravity. In our theory, the disformal frame is the natural choice for the minimal coupling of the matter to gravity which avoids the runaway modes as well as the strong coupling of the tensor modes.
@@ 6 Conclusion
In this work, we have studied Proca theory, non-minimally coupled to gravity, with two types of couplings – one with the Ricci scalar and the Ricci tensor. The cosmological aspects of this theory have recently drawn a lot of attention, making it an intriguing possibility to underline the gravitational production of dark matter and a possible candidate for generating primordial magnetic fields and inflationary candidates.
By studying this theory in flat space-time, with metric perturbations also taken into account, we have found a somewhat surprising result when considering the massless limit. On the one hand, the interaction of the vector field with the Ricci scalar results in the strong coupling of the longitudinal modes. This is natural, as the longitudinal modes are absent in the massless theory. Thus, similarly to the Proca theory in the presence of quartic and cubic self-interactions, the longitudinal modes enter a strong coupling regime once the corresponding non-linear term becomes of the same order as the linear one. Beyond the strong coupling scale, the vector and tensor modes remain weakly coupled, up to the Planck length, due to the amplitude of quantum fluctuations for the tensor modes.
@@
The coupling with the Ricci tensor, on the other hand, indicates that both tensor and longitudinal modes become strongly coupled at the same length scale. This, however, is an unphysical result – it would imply that if one allows for such non-minimal coupling in nature, and a photon that has a small mass, the gravitational waves would quickly lose their linear propagator. Curiously, this is not the only complication that arises when the coupling with the Ricci tensor is taken into account. A recent study has found a new class of negative frequency modes in curved space-time, known as the runaway modes in the presence of non-minimal coupling [26]. In this work, we have shown that the origin of such modes lies in the coupling of the vector field to the Ricci tensor, which is problematic even in the flat background. One should note that the coupling with the Ricci scalar, on the other hand, does not give rise to such modes.
@@
Nevertheless, by carefully analyzing the theory at this strong coupling scale, we have found that the tensor modes have to be transformed. A similar scenario was found also in massive Yang-Mills theory, with mass added by hand, where with standard decomposition, the transverse modes were becoming strongly coupled. However, by correcting their definition, it was shown that ultimately, these modes remain weakly coupled at all scales. In the case of non-minimally coupled Proca theory, we have found that the introduction of corresponding disformal coupling in the theory simultaneously removes the inconsistent interaction by introducing the necessary coupling. This leads us to the self-interaction of the longitudinal modes at the quartic order, becoming a necessary step to make the theory safe. Curiously, in this disformal frame, the strong coupling scale is left unchanged, in contrast to the case of mYM theory.
Notably, in flat spacetime, the form of the disformal coupling can also be deducted by directly studying the coupling with the Ricci tensor and the vector fields on the level of action. In this case, the Ricci tensor depends only on the perturbations, so we can trace the appearance of the interaction between the tensor modes and the longitudinal dof. In the cosmological backgrounds, this is no longer apparent, as the background value of the Ricci tensor is non-vanishing. Curiously, however, the disformal frame is defined by the same form of the disformal transformation. In it, the runaway issue is absent due to the disformal transformation which at the same generates further self-interactions of the vector field. It is important to stress, that in this frame the matter is minimally coupled to gravity, which guarantees that the runaway modes will not re-appear in a more complicated way. Given that, in this frame, the instability is absent (and the remnant of the coupling between the Ricci tensor remains in the higher couplings), and at the same time it resolves the strong coupling of the tensor modes in the Minkowski background, it indicates that the theory formulated in this way will be healthy also for general background.
While the disformal frame is one possible solution to the runaway and the strong coupling of tensor modes, one might regard it as fine-tuning. In a more general setting, the instability is expected to remain. For instance, if the matter is minimally coupled to gravity in the original frame, the runaway might reappear in a non-trivial way. It would be nevertheless interesting to study this case and confirm if the instability persists, or vanishes as well.
Another interesting aspect is to study the quantized theory and to explore further aspects of the strong coupling in curved space-time for this type of theory. In particular, it would be intriguing to see if such a possibility could resolve the well-known ghost, that appears for $M_{T}^{2}<0$. Finally, one should note that while in Proca theory, the longitudinal modes are the source of the singularities in mass within the perturbative regime, in the Kalb-Ramond field, this role is taken by the vector modes [50]. Therefore, it will be interesting to also extend this analysis to this case, and study its consequences.